OpenStudy (anonymous):

How to solve 2X^2-4X-4=0 by completing the square. Please explain thoroughly!

OpenStudy (anonymous):

first you must divide the whole equation by the leading coefficient (2) so you get x^2-2x-2=0 then you can add the 2 to the other side x^2-2x=2 then you take "b" (-2 from -2x in this case) and do (b/2)^2 (-2/2)^2=1 now we can add 1 to the equation IF we add it to both sides x^2-2x+1=2+1 Now you can factor the left side and then subtract the 3 over to the left. Does that help?

OpenStudy (anonymous):

Is it X=1+/- Root(3)

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Thanks! I hope its right tho; it is for a project!